OpenStudy (anonymous):

What is the length of the hypotenuse of a right triangle whose legs are 5 and √2?

OpenStudy (anonymous):

Not one bit .-. I have no idea what that question is even asking

OpenStudy (anonymous):

 You're finding x

OpenStudy (anonymous):

Idk how...

OpenStudy (anonymous):

\[x^2+y^2 = c^2\] \[\sqrt{2}^2 +5^2 = ?\]

OpenStudy (anonymous):

27?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

So is it 2√3 or 3√3 or √29?

OpenStudy (kagıtucak):

\[\sqrt{27}=\sqrt{9.3}=3\sqrt{3}\]

OpenStudy (anonymous):

Thank you ^-^